OpenStudy (anonymous):

WILL GIVE MEDAL PLEASE HELP!!! The container that holds the water for the football team is 1/10 full. After pouring in 10 gallons of water, it is 1/2 full. How many gallons can the container hold?

OpenStudy (anonymous):

hold on let me do the math real quick

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

im almost positive that its 9/10

OpenStudy (anonymous):

was i right?

OpenStudy (anonymous):

It doesn't give the option to do a fraction... :(

OpenStudy (anonymous):

And 9 isn't correct

OpenStudy (anonymous):

what options do you have

OpenStudy (anonymous):

is it multiple choice or nah

OpenStudy (anonymous):

It doesn't give me options I have to write it in myself

OpenStudy (anonymous):

Yeah its not :(

OpenStudy (anonymous):

The answer was 25 gallons...

OpenStudy (anonymous):

20?

OpenStudy (anonymous):

Kala's penny bank is 2/3 full. After she removes 80 pennies, it is 1/2 full. How many pennies can Kala's bank hold?

OpenStudy (phi):

call the unknown amount of gallons in a full bucket "x" they say 1/10 full 1/10 full is 1/10 of x or 1/10 * x or x/10 ok so far?

OpenStudy (anonymous):

I dont need the first answer anymore cause it timed out

OpenStudy (phi):

do you want to do the penny problem ?

OpenStudy (anonymous):

Yes!

OpenStudy (phi):

How many pennies can Kala's bank hold? we want to know how many pennies. Call that amount "x" do you know how to write 2/3 of that amount ?

OpenStudy (anonymous):

2/3 of x? I don't have the number how can I write 2/3 of it

OpenStudy (phi):

Kala's penny bank is 2/3 full. you want to write 2/3 of what ? (what letter do we use?) yes you write 2/3 of x which means 2/3 times x or just \[ \frac{2}{3} x \]

OpenStudy (anonymous):

This is why math makes absolutely no sense to me sorry haha I just want to know the formula so I can do the other problems

OpenStudy (anonymous):

OH okay

OpenStudy (phi):

It is a way of thinking, and the only reason to do these is to learn how to think this way. It's like learning to read... hard until you "get it", then it's useful. anyway she removes 80 pennies, that means you have 2/3 x and you take away 80 how do you write "take away 80" hint: add or subtract 80?

OpenStudy (anonymous):

Subtract

OpenStudy (phi):

if it is not clear, how would you do: you have 100 and take away 80 how do you write that problem ?

OpenStudy (phi):

you have 100 and take away 80 you would write 100 - 80 your problem is you have 2/3 x and take away 80 how do you write that ?

OpenStudy (phi):

type faster or it will time out

OpenStudy (phi):

2/3 x - 80

OpenStudy (phi):

bank is 2/3 full. After she removes 80 pennies, it is 1/2 full. that says after you subtract 80 you have 1/2 full. How do you write 1/2 of "full" you write 1/2 x so the problem is 2/3 x - 80 = 1/2 x to solve this, add -1/2 x to both sides 2/3 x - 1/2 x -80 = 1/2 x - 1/2 x now simplify the right side to zero 2/3 x - 1/2 x -80=0 now add +80 to both sides can you do that ?

OpenStudy (anonymous):

Sorry had to go in the shower, there's no way to just plug in the numbers to an equation and figure it out from there?? Is all this work really necessary?

OpenStudy (phi):

yes, this is how you do this problem.